# When Geoscience Meets Foundation Models: Towards General Geoscience Artificial Intelligence System

### v4

@@ Abstract
Artificial intelligence (AI) has significantly advanced Earth sciences, yet its full potential in to comprehensively modeling Earth’s complex dynamics remains unrealized. Geoscience foundation models (GFMs) emerge as a paradigm-shifting solution, integrating extensive cross-disciplinary data to enhance the simulation and understanding of Earth system dynamics. These data-centric AI models extract insights from petabytes of structured and unstructured data, effectively addressing the complexities of Earth systems that traditional models struggle to capture. The unique strengths of GFMs include flexible task specification, diverse input-output capabilities, and multi-modal knowledge representation, enabling analyses that surpass those of individual data sources or traditional AI methods. This review not only highlights the key advantages of GFMs, but also presents essential techniques for their construction, with a focus on transformers, pre-training, and adaptation strategies. Subsequently, we examine recent advancements in GFMs, including large language models, vision models, and vision-language models, particularly emphasizing the potential applications in remote sensing. Additionally, the review concludes with a comprehensive analysis of the challenges and future trends in GFMs, addressing five critical aspects: data integration, model complexity, uncertainty quantification, interdisciplinary collaboration, and concerns related to privacy, trust, and security. This review offers a comprehensive overview of emerging geoscientific research paradigms, emphasizing the untapped opportunities at the intersection of advanced AI techniques and geoscience. It examines major methodologies, showcases advances in large-scale models, and discusses the challenges and prospects that will shape the future landscape of GFMs. The paper highlights a dynamic field rich with possibilities, poised to unlock new insights into Earth’s complexities and further advance geoscience exploration.
@@
Artificial intelligence, geoscience foundation models (GFMs), large models, remote sensing, Earth system science.
@@ I Introduction
[lines=2]AS Earth science enters the era of big data, artificial intelligence (AI) holds significant potential not only for addressing classic geoscientific problems and enhancing understanding of Earth’s complex, interactive, and multiscale processes [1, 2]. The availability of massive volumes of Earth system data, alread exceeding dozens of petabytes with hundreds of terabytes transmitted daily, has facilitated the widespread adoption of AI technologies, in data-driven Earth system sciences [3, 4]. The utilization of AI not only enhances geological events (e.g., earthquakes and volcanic eruptions) through real-time seismic data analysis but also holds significant value in bias correction and upsampling of Earth system models [5, 1, 6, 7, 8]. Machine learning approaches support the exploration and mapping of natural resources (e.g., minerals and oil), thereby increasing accuracy and minimizing the ecological impacts of mining and drilling operations [9, 10, 11]. Moreover, the application of deep learning algorithms in climate modeling and weather forecasting plays a crucial role in understanding the effects of climate change, refining weather predictions, and enhancing disaster management and preparedness [12, 13, 14, 15, 16, 17]. Additionally, the advanced capabilities of AI in processing and interpreting extensive amounts of satellite imagery and sensor data also makes a substantial contribution to environmental monitoring and conservation, which enables a diverse array of critical applications, including environmental monitoring, land cover classification, urban planning, disaster response, and climate change analysis [18, 19, 20, 21].
@@
Although there have been preliminary efforts to develop geoscience foundation models (GFMs) [47, 48, 8, 49, 45, 50, 46], widespread adoption within geoscience AI remains limited. The slow progression in GFMs is attributed to persistent challenges in accessing and processing multi-source and multi-modal geoscience datasets, the inherent complexity of Earth systems, and the inherent limitations of current AI technologies [51, 52, 53]. Although the task-specific approach persists as the dominant paradigm in geoscience AI. These innovations in GFMs may transform how geoscientists process and interpret data, leading to enhanced accuracy and impact in critical research areas like sustainability, climate change, extreme weather events, natural hazards, and environmental monitoring [54, 55, 56]. Therefore, the adoption of adaptable, multifaceted foundation models in place of narrow, static task-specific models has the potential to not only provide new insights and capabilities to Earth science, but also to transform our understanding of complex Earth systems, drive innovation in interdisciplinary research, and address global challenges with greater precision and foresight.
@@ I-A Existing Surveys and Reviews
Table I provides an overview of existing surveys and reviews related to foundation models for geoscience. These works have explored various aspects of foundation models in geoscience, including remote sensing, urban science, vision-language models, generative artificial intelligence, geophysical foundation models, and geospatial location embedding. These works are published in 2023 and 2024. For example, Jiao et al. [57] contributed to the field with their proposed brain-inspired framework for remote sensing foundation models. In a Chinese-language publication, Zhang’s team [58] offered a comprehensive review of pre-training techniques for large remote sensing models. The landscape of smart cities is enriched by the work’s [55] introduction of Urban Foundation Models. The authors in [45] delved into the intersection of vision-language models and remote sensing in their examination. The potential of generative AI and large language models in geoscience is explored in depth by the research group led by Hadid et al. [59]. A thought-provoking discussion on geophysical foundation models in exploration geophysics was presented in the perspective piece by the work [60]. Zhou et al. spearheaded a thorough review [61] that sheds light on vision-language geo-foundation models. The authors in [62] break new ground with their investigation into geospatial location embedding in large language models. However, there is still a need for a comprehensive survey that integrates these diverse perspectives and provides a holistic view of the current state of the field. This paper aims to fill this gap by providing a comprehensive review of the key advantages, techniques, and advances in GFMs. We also discuss the challenges and future trends of GFMs in geoscience, highlighting the potential of these models to revolutionize Earth system science.
@@
2024 This work A comprehensive review of geoscience foundation models (GFMs) from the view of key advantages, key techniques, and key advances.
@@ I-B Contributions and Organization
  • This review provides a comprehensive overview of geoscience foundation models (GFMs) from the perspective of key advantages, key techniques, and key advances. The study also discusses the challenges and future trends of GFMs in geoscience, emphasizing the potential of these models to revolutionize Earth system science.
@@
  • We present a comprehensive review of the key advances in GFMs, including large language models (LLMs), large vision models (LVMs), and large vision-language models (LVLMs), with a particular emphasis on their applications in remote sensing. We also discuss the potential of GFMs to revolutionize Earth system science by providing deep, actionable insights into complex Earth systems.
@@
The remainder of this paper is organized as follows. Section II discusses the key advantages of GFMs, including their data-centric paradigm, scalability, generalizability, dynamic interactions, and causal inference. Section III introduces the geoscience domain knowledge and the potential applications of GFMs in Earth science. Section IV presents the recent advances in GFMs, including large language models, large vision models, and large vision-language models, especially remote sensing applications. Section V discusses the challenges and future trends of GFMs. Finally, Section VI concludes the paper.
@@ II Key Advantages of Geoscience Foundation Models
In this section, we discuss the key advantages of geoscience foundation models (GFMs) from the view of a new paradigm, Geoscience domain knowledge, and application domains. Data-centric AI, ccalability and generalizability, and dynamic interactions make GFMs a new paradigm compared to traditional AI. Geoscience domain knowledge including publiocations, geological data, and knowledge graph further enhance the capabilities of GFMs. The application domain of GFMs in Earth science are also discussed, including space, air, ground, and ocean domains.
@@ II-A2 Scalability and Generalizability
Foundation models exhibit strong generalizability due to their training methodology. Specifically, foundation models are pretrained on diverse datasets in a self-supervised manner, learning broad representations of language, vision, robotics, or other modalities. This differs from traditional supervised learning, where models are tailored to specific labeled datasets. The self-supervised pretraining endows foundation models with extensive world knowledge and versatility. Subsequently, the pretrained models can be adapted and fine-tuned for myriad downstream tasks using relatively little labeled data. Even unseen tasks can be potentially tackled in a zero-shot setting without any downstream data. Their flexibility and wide applicability highlight the generalizability of foundation models across domains. Their generalizability originates from the self-supervised pretraining and enables adaption to new distributions and tasks.
@@ II-A3 Dynamic Interactions
GFMs aim to represent complex natural processes in Earth system like mantle convection, crustal deformation, groundwater flow, ocean and atmospheric dynamics, or vegetation. These processes interact in non-linear ways across multiple spatial and temporal scales. GFMs can be taught to flexibly specify new tasks without requiring models to be retrained. For instance, a weather prediction model can be realized by training on historical climate data, and this model can also be helpful for the forecast of rainfall-induced shallow landslides.
@@
The integration of GFMs into the exploration of Earth’s processes and phenomena heralds a transformative shift in our ability to discern and understand the intricate causal relationships inherent in geoscience knowledge [65]. By leveraging data from diverse sources and modalities—ranging from satellite imagery and sensor data to historical climate records and geological samples—GFMs offer a novel approach to modeling the Earth system that is both holistic and nuanced. This comprehensive data integration enables GFMs to capture the complexity and interconnectedness of the Earth system’s components. For instance, by analyzing patterns across atmospheric data, oceanic temperatures, and land use changes, GFMs can help researchers identify underlying causal mechanisms driving climate change, such as the feedback loops between greenhouse gas emissions and global warming [66, 67]. Similarly, in the context of natural disasters, GFMs can elucidate the cascading effects of a single event, like how an earthquake can trigger tsunamis, which in turn affect coastal erosion and sediment transport. The potential of GFMs to elucidate causal relationships in geoscience is grounded in their ability to process and analyze vast datasets that would be unmanageable for traditional analytical methods. Through advanced machine learning techniques, GFMs can detect patterns and correlations within this data, proposing hypotheses about causal links that can be further tested through empirical research or simulation models. This not only accelerates the discovery of new knowledge but also enhances our ability to predict future Earth system changes and inform policy and decision-making. Moreover, the flexibility of GFMs to incorporate data from new sources and update their models in real-time allows for the continuous refinement of our understanding of causal relationships as new evidence emerges. This adaptability is crucial in geoscience, where ongoing changes in the Earth system and technological advancements in data collection methods constantly evolve the landscape of available information.
@@ II-B Geoscience Domain Knowledge
As shown in Fig. 2, geoscience encompasses diverse domains spanning the physical makeup and processes of the Earth. Core geoscience domains include geology, geography, geophysics, hydrology, oceanography, meteorology, and planetary science. Each domain has accumulated extensive observations, data, and models of a wide range of Earth system components. For instance, geology involves studying the composition, structure, and history of the Earth’s crust and interior, accumulating knowledge on rock types, plate tectonics, mineral resources, and geologic time [68]. Meanwhile, meteorology focuses on the atmosphere and atmospheric phenomena like weather patterns, clouds, and precipitation. Key meteorological knowledge includes air masses, fronts, forecasting models, and climate dynamics [69]. Moreover, geography tends to explore the spatial patterns and processes of natural phenomena, such as climate change, natural disasters, and biodiversity, as well as human aspects like urbanization, cultural landscapes, and economic development. Geographers require a diverse set of knowledge, including understanding physical geography, human geography, and technical skills in geographic information systems (GIS), remote sensing, and spatial analysis [70].
@@
Figure 2: Overview of the Geoscience foundation model (GFM). By leveraging techniques like self-supervised learning, the GFMs undergo training using multimodal and multidimensional geoscience data models. For enhanced versatility, the data schema can be coupled with language inputs, such as text, voice, or video data. Furthermore, the GFMs require access to diverse sources of geoscientific knowledge to facilitate geoscientific inference tasks, unlocking a plethora of functionalities applicable to downstream applications. Once trained, the GFMs can dynamically perform user-specified tasks in real time. To accomplish this, the model can retrieve contextual information from geological mapping manuscripts, knowledge graphs, databases, and other relevant sources, capitalizing on formal geoscience knowledge to deduce solutions for previously unseen tasks.
II-C Application Domain
@@ II-C3 Ground
II-C4 Ocean
AI models enable holistic monitoring, modeling, and forecasting of conditions in the near-surface ocean through multi-sensor data analysis, digital twin simulations, and long-term observations at multi-scale [76]. However, compared with the near-surface ocean, the intermediate ocean at a depth from 100 to 1000 m is confronted with a series of challenges primarily stemming from the scarceness of observational data, which has emerged as a significant impediment constraining the progressive evolution of marine science and data-driven modelling in oceanography. We anticipate the development of a GFM featuring real-time, immersive, and highly sensitive dynamic interactive visualization capabilities for three-dimensional oceanic data. This model holds the promise of gradually unraveling the enigma surrounding the intermediate and deep oceans, thereby ushering future marine science into a transformative new phase of exploration and understanding. For instance, advanced GFMs have the potential to catalyze a transformation in midocean remote sensing technology, offering the capability for autonomous, long-term monitoring of poorly explored mid-ocean regions. This can be achieved through intelligent fusion and analysis of diverse subsurface, environmental, and biological data streams from distributed, next-generation ocean observation platforms, thereby revolutionizing the way mid-ocean regions are studied and understood. Moreover, GFMs play a transformative role within the framework of Deep Blue AI [77], facilitating advanced data analysis, predictive modeling, and well-informed decision-making to elevate the quality of marine science research, exploration, and conservation endeavors, thereby helping to unraveling the enigma of one of Earth’s least-explored yet profoundly significant frontiers.
III Key Techniques of Constructing Geoscience Foundation Models

### v5

@@ Abstract
Artificial intelligence (AI) has significantly advanced Earth sciences, yet its full potential in to comprehensively modeling Earth’s complex dynamics remains unrealized. Geoscience foundation models (GFMs) emerge as a paradigm-shifting solution, integrating extensive cross-disciplinary data to enhance the simulation and understanding of Earth system dynamics. These data-centric AI models extract insights from petabytes of structured and unstructured data, effectively addressing the complexities of Earth systems that traditional models struggle to capture. The unique strengths of GFMs include flexible task specification, diverse input-output capabilities, and multi-modal knowledge representation, enabling analyses that surpass those of individual data sources or traditional AI methods. This review not only highlights the key advantages of GFMs, but also presents essential techniques for their construction, with a focus on transformers, pre-training, and adaptation strategies. Subsequently, we examine recent advancements in GFMs, including large language models, vision models, vision-language models, and foundation-model-based agents, particularly emphasizing the potential applications in remote sensing. Additionally, the review concludes with a comprehensive analysis of the challenges and future trends in GFMs, addressing five critical aspects: data integration, model complexity, uncertainty quantification, interdisciplinary collaboration, and concerns related to privacy, trust, and security. This review offers a comprehensive overview of emerging geoscientific research paradigms, emphasizing the untapped opportunities at the intersection of advanced AI techniques and geoscience. It examines major methodologies, showcases advances in large-scale models, and discusses the challenges and prospects that will shape the future landscape of GFMs. The paper highlights a dynamic field rich with possibilities, poised to unlock new insights into Earth’s complexities and further advance geoscience exploration.
@@
Artificial intelligence, geoscience foundation models (GFMs), foundation models, remote sensing, Earth system science.
@@ I Introduction
[lines=2]AS Earth science enters the era of big data, artificial intelligence (AI) holds significant potential not only for addressing classic geoscientific problems and enhancing understanding of Earth’s complex, interactive, and multiscale processes [1, 2]. The availability of massive volumes of Earth system data, already exceeding dozens of petabytes with hundreds of terabytes transmitted daily, has facilitated the widespread adoption of AI technologies, in data-driven Earth system sciences [3, 4]. The utilization of AI not only enhances geological events (e.g., earthquakes and volcanic eruptions) through real-time seismic data analysis but also holds significant value in bias correction and upsampling of Earth system models [5, 1, 6, 7, 8]. Machine learning approaches support the exploration and mapping of natural resources (e.g., minerals and oil), thereby increasing accuracy and minimizing the ecological impacts of mining and drilling operations [9, 10, 11]. Moreover, the application of deep learning algorithms in climate modeling and weather forecasting plays a crucial role in understanding the effects of climate change, refining weather predictions, and enhancing disaster management and preparedness [12, 13, 14, 15, 16, 17]. Additionally, the advanced capabilities of AI in processing and interpreting extensive amounts of satellite imagery and sensor data also makes a substantial contribution to environmental monitoring and conservation, which enables a diverse array of critical applications, including environmental monitoring, land cover classification, urban planning, disaster response, and climate change analysis [18, 19, 20, 21].
@@
Although there have been preliminary efforts to develop geoscience foundation models (GFMs) [47, 48, 8, 49, 45, 50, 46], widespread adoption within geoscience AI remains limited. The slow progression in GFMs is attributed to persistent challenges in accessing and processing multi-source and multi-modal geoscience datasets, the inherent complexity of Earth systems, and the inherent limitations of current AI technologies [51, 52, 53]. Although the task-specific approach persists as the dominant paradigm in geoscience AI. These innovations in GFMs may transform how geoscientists process and interpret data, leading to enhanced accuracy and impact in critical research areas such as sustainability, climate change, extreme weather events, natural hazards, and environmental monitoring [54, 55, 56]. Therefore, the adoption of adaptable, multifaceted foundation models in place of narrow, static task-specific models has the potential to not only provide new insights and capabilities to Earth science, but also to transform our understanding of complex Earth systems, drive innovation in interdisciplinary research, and address global challenges with greater precision and foresight.
@@ I-A Existing Surveys and Reviews
Table I provides an overview of existing surveys and reviews related to foundation models for geoscience. These works have explored various aspects of foundation models in geoscience, including remote sensing, urban science, vision-language models, generative artificial intelligence, geophysical foundation models, and geospatial location embedding. These works are published in 2023 and 2024. For example, Jiao et al. [57] contributed to the field with their proposed brain-inspired framework for remote sensing foundation models. In a Chinese-language publication, Zhang’s team [58] offered a comprehensive review of pre-training techniques for large remote sensing models. The landscape of smart cities is enriched by the work’s [55] introduction of Urban Foundation Models. The authors in [45] delved into the intersection of vision-language models and remote sensing in their examination. The potential of generative AI and large language models in geoscience is explored in depth by the research group led by Hadid et al. [59]. A thought-provoking discussion on geophysical foundation models in exploration geophysics was presented in the perspective piece by the work [60]. Zhou et al. spearheaded a thorough review [61] that sheds light on vision-language geo-foundation models. The authors in [62] break new ground with their investigation into geospatial location embedding in large language models. However, there is still a need for a comprehensive survey that integrates these diverse perspectives and provides a holistic view of the current state of the field. This paper aims to fill this gap by providing a comprehensive review of the key advantages, applications, techniques, and advances in GFMs. We also discuss the challenges and future trends of GFMs in geoscience, highlighting the potential of these models to revolutionize Earth system science.
@@
2024 This work A comprehensive review of geoscience foundation models (GFMs) from the view of key advantages, key applications, key techniques, and key advances.
@@ I-B Contributions and Organization
  • This review provides a comprehensive overview of geoscience foundation models (GFMs) from the perspective of key advantages, key applications, key techniques, and key advances. The study also discusses the challenges and future trends of GFMs in geoscience, emphasizing the potential of these models to revolutionize Earth system science.
@@
  • We present a comprehensive review of the key advances in GFMs, including large language models (LLMs), large vision models (LVMs), large vision-language models (LVLMs), and foundation-model-based agents, with a particular emphasis on their applications in remote sensing. We also discuss the potential of GFMs to revolutionize Earth system science by providing deep, actionable insights into complex Earth systems.
@@
The remainder of this paper is organized as follows. Section II discusses the key advantages of GFMs, including their data-centric paradigm, scalability, generalizability, dynamic interactions, and causal inference. Section III explores the potential applications of GFMs in Earth science, including space, air, ground, and ocean domains. Section IV introduces the geoscience domain knowledge and the potential applications of GFMs in Earth science. Section V presents the recent advances in GFMs, including large language models, large vision models, large vision-language models, and foundation-model-based agents, especially remote sensing applications. Section VI discusses the challenges and future trends of GFMs. Finally, Section VII concludes the paper.
@@ II Key Advantages of Geoscience Foundation Models
In this section, we discuss the key advantages of geoscience foundation models (GFMs) from the view of a new paradigm, Geoscience domain knowledge, and application domains. Data-centric AI, capability and generalizability, and dynamic interactions make GFMs a new paradigm compared to traditional AI. Geoscience domain knowledge including publications, geological data, and knowledge graph further enhance the capabilities of GFMs. The application domain of GFMs in Earth science are also discussed, including space, air, ground, and ocean domains.
@@ II-A2 Scalability and Generalizability
Foundation models exhibit strong generalizability due to their training methodology. Specifically, foundation models are pre-trained on diverse datasets in a self-supervised manner, learning broad representations of language, vision, robotics, or other modalities. This differs from traditional supervised learning, where models are tailored to specific labeled datasets. The self-supervised pre-training endows foundation models with extensive world knowledge and versatility. Subsequently, the pretrained models can be adapted and fine-tuned for myriad downstream tasks using relatively little labeled data. Even unseen tasks can be potentially tackled in a zero-shot setting without any downstream data. Their flexibility and wide applicability highlight the generalizability of foundation models across domains. Their generalizability originates from the self-supervised pre-training and enables adaption to new distributions and tasks.
@@ II-A3 Dynamic Interactions
GFMs aim to represent complex natural processes in Earth system such as mantle convection, crustal deformation, groundwater flow, ocean and atmospheric dynamics, or vegetation. These processes interact in non-linear ways across multiple spatial and temporal scales. GFMs can be taught to flexibly specify new tasks without requiring models to be retrained. For instance, a weather prediction model can be realized by training on historical climate data, and this model can also be helpful for the forecast of rainfall-induced shallow landslides.
@@
The integration of GFMs into the exploration of Earth’s processes and phenomena heralds a transformative shift in our ability to discern and understand the intricate causal relationships inherent in geoscience knowledge [65]. By leveraging data from diverse sources and modalities—ranging from satellite imagery and sensor data to historical climate records and geological samples—GFMs offer a novel approach to modeling the Earth system that is both holistic and nuanced. This comprehensive data integration enables GFMs to capture the complexity and interconnectedness of the Earth system’s components. For instance, by analyzing patterns across atmospheric data, oceanic temperatures, and land use changes, GFMs can help researchers identify underlying causal mechanisms driving climate change, such as the feedback loops between greenhouse gas emissions and global warming [66, 67]. Similarly, in the context of natural disasters, GFMs can elucidate the cascading effects of a single event, such as how an earthquake can trigger tsunamis, which in turn affect coastal erosion and sediment transport. The potential of GFMs to elucidate causal relationships in geoscience is grounded in their ability to process and analyze vast datasets that would be unmanageable for traditional analytical methods. Through advanced machine learning techniques, GFMs can detect patterns and correlations within this data, proposing hypotheses about causal links that can be further tested through empirical research or simulation models. This not only accelerates the discovery of new knowledge but also enhances our ability to predict future Earth system changes and inform policy and decision-making. Moreover, the flexibility of GFMs to incorporate data from new sources and update their models in real-time allows for the continuous refinement of our understanding of causal relationships as new evidence emerges. This adaptability is crucial in geoscience, where ongoing changes in the Earth system and technological advancements in data collection methods constantly evolve the landscape of available information.
@@ II-B Geoscience Domain Knowledge
As shown in Fig. 2, geoscience encompasses diverse domains spanning the physical makeup and processes of the Earth. Core geoscience domains include geology, geography, geophysics, hydrology, oceanography, meteorology, and planetary science. Each domain has accumulated extensive observations, data, and models of a wide range of Earth system components. For instance, geology involves studying the composition, structure, and history of the Earth’s crust and interior, accumulating knowledge on rock types, plate tectonics, mineral resources, and geologic time [68]. Meanwhile, meteorology focuses on the atmosphere and atmospheric phenomena such as weather patterns, clouds, and precipitation. Key meteorological knowledge includes air masses, fronts, forecasting models, and climate dynamics [69]. Moreover, geography tends to explore the spatial patterns and processes of natural phenomena, such as climate change, natural disasters, and biodiversity, as well as human aspects such as urbanization, cultural landscapes, and economic development. Geographers require a diverse set of knowledge, including understanding physical geography, human geography, and technical skills in geographic information systems (GIS), remote sensing, and spatial analysis [70].
@@
Figure 2: Overview of the Geoscience foundation model (GFM). By leveraging techniques such as self-supervised learning, the GFMs undergo training using multimodal and multidimensional geoscience data models. For enhanced versatility, the data schema can be coupled with language inputs, such as text, voice, or video data. Furthermore, the GFMs require access to diverse sources of geoscientific knowledge to facilitate geoscientific inference tasks, unlocking a plethora of functionalities applicable to downstream applications. Once trained, the GFMs can dynamically perform user-specified tasks in real time. To accomplish this, the model can retrieve contextual information from geological mapping manuscripts, knowledge graphs, databases, and other relevant sources, capitalizing on formal geoscience knowledge to deduce solutions for previously unseen tasks.
III Key Applications of Geoscience Foundation Models
@@ III-C Ground
III-D Ocean
AI models enable holistic monitoring, modeling, and forecasting of conditions in the near-surface ocean through multi-sensor data analysis, digital twin simulations, and long-term observations at multi-scale [76]. However, compared with the near-surface ocean, the intermediate ocean at a depth from 100 to 1000 m is confronted with a series of challenges primarily stemming from the scarceness of observational data, which has emerged as a significant impediment constraining the progressive evolution of marine science and data-driven modelling in oceanography. We anticipate the development of a GFM featuring real-time, immersive, and highly sensitive dynamic interactive visualization capabilities for three-dimensional oceanic data. This model holds the promise of gradually unraveling the enigma surrounding the intermediate and deep oceans, thereby ushering future marine science into a transformative new phase of exploration and understanding. For instance, advanced GFMs have the potential to catalyze a transformation in mid-ocean remote sensing technology, offering the capability for autonomous, long-term monitoring of poorly explored mid-ocean regions. This can be achieved through intelligent fusion and analysis of diverse subsurface, environmental, and biological data streams from distributed, next-generation ocean observation platforms, thereby revolutionizing the way mid-ocean regions are studied and understood. Moreover, GFMs play a transformative role within the framework of Deep Blue AI [77], facilitating advanced data analysis, predictive modeling, and well-informed decision-making to elevate the quality of marine science research, exploration, and conservation endeavors, thereby helping to unraveling the enigma of one of Earth’s least-explored yet profoundly significant frontiers.
IV Key Techniques of Constructing Geoscience Foundation Models